The length of a rectangle is 9 1/2 inches and the width is 2 3/4 inches. What is the ratio, using whole numbers, of the length to the width? I need all the work shown. I know the answer, but not how to solve it. Thanks so much!
Answer by stanbon(75887)

The length of a rectangle is 9 1/2 inches and the width is 2 3/4 inches. What is the ratio, using whole numbers, of the length to the width?
-------
9 1/2 = 19/2
------
2 3/4 = 11/4
-----
(19/2)/(11/4) = 38/11
----
Ratio:: 38:11
======================
Cheers,
Stan H.